OpenStudy (anonymous):

Determine the dimensions of a rectangular solid (with a square base) of maximum volume if its surface area is 150 square inches. base dimension

OpenStudy (anonymous):

I NEED HELP I THINK THE ANSWER IS 5*5*5 BUT EVERYTIME I SUBMIT THAT IT DOES NOT WORK!!

OpenStudy (anonymous):

I got exactly the same yours. but I think may be the | 5 | will work because when solving for x ( the side of base) we get x^2 =25 . In fact the length of side must be positive and everybody knows that except the computer. try again.

OpenStudy (anonymous):



OpenStudy (anonymous):

\[V(x,y,z)=xyz\\ S(x,y,z)=2(xy)+2(yz)+2(zy)=150 \]

OpenStudy (anonymous):

the problem gives us " square base", does it mean y =x?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

\[V(x,y)=x^2y\\S(x,y)=2x^2+2xy+2xy=150\implies x(x+2y)=75\]

OpenStudy (anonymous):

now, the cost function becomes, \[V(x)=x^2{1\over2}\left({75\over x}-x\right)={1\over2}(75x-x^3)\] maximize this the ususal way

OpenStudy (anonymous):

exactly 5,5,5

OpenStudy (anonymous):

but he said that it doesn't work when he submited

OpenStudy (anonymous):

i am submitting it as 5x5x5 and i tried 5*5*5 and it does not work...

OpenStudy (anonymous):

thank you both for your help i really appreciate it, i just dont know why it is marking it wrong? should it be written in a different way?

OpenStudy (anonymous):

\[V'(x)={75\over2}-{3\over2}x^2=0\implies x^2-15=0\implies x=\pm\sqrt{15}\]

OpenStudy (anonymous):

that does not work either.

OpenStudy (anonymous):

:) length of a side cannot be negative. hope you put in the positive squre-root

OpenStudy (anonymous):

i cannot submit it as square root it does not give me that option..

OpenStudy (anonymous):

so, should be 3.8 or somethign like that

OpenStudy (anonymous):

hi friend, how can you get x =+- sqr 15? i don't know

OpenStudy (anonymous):

hehe. HINT: 15 lies between what two perfect squares? -> 9 and 16

OpenStudy (anonymous):

and then close to which one

OpenStudy (anonymous):

why don't we go on the simplest way to get x ?

OpenStudy (anonymous):

@Hoa I do not see what you mean

OpenStudy (anonymous):

\[\frac{ 75 }{ 2 }-\frac{ 3x^2 }{ 2 }=0 --->\frac{ 3x^2 }{ 2 }= \frac{ 75 }{ 2 }---> x^2=25\]

OpenStudy (anonymous):

ooohh "25".. I put "15".. duh

OpenStudy (anonymous):

hihihi.... magic mathematic ?

OpenStudy (anonymous):

@hayatabz1 just to verify, is the solid closed or open?

OpenStudy (anonymous):

like there is no suface on the top!!!

OpenStudy (anonymous):

yeah... I know what you mean

OpenStudy (anonymous):

but i dont' think it is the case.

OpenStudy (anonymous):

it is asking for "base dimension". I do not think it includes the height. Those online classes are so crazy sometomes. Esp with physics and numbers

OpenStudy (anonymous):

you mean it's vector of base?

OpenStudy (anonymous):

so, it's \[\left(\begin{matrix}1 \\ 1\\1\end{matrix}\right)\]

OpenStudy (anonymous):

cant say. dont have "O". but it says base dimensions..

OpenStudy (anonymous):

I surrender!!! cannot follow anymore. too complicated.. if it's a vector, I'm ok. if it's open solid, it engages to chemistry. I am hand off, long time ago when I studied it.

OpenStudy (anonymous):

no. just subtract one of the surfaces from S(x,y)

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

ok. let the asker does his work. I think after 3 times submit, he has no more option for re submit the answer. he lost it

OpenStudy (aravindg):

Hi there! Welcome to Openstudy :) If you have any doubts on how the site works ,feel free to ask me or any other Ambassador with purple A near them .Have great time here . Check out our COC : http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

its a solid

OpenStudy (anonymous):

hold on, in the question does it say "square base" or is that somethig you came up with?

OpenStudy (anonymous):

it says exactly this..

OpenStudy (anonymous):

and im looking for the base dimension

OpenStudy (anonymous):

Determine the dimensions of a rectangular solid (with a square base) of maximum volume if its surface area is 150 square inches. base dimension

OpenStudy (anonymous):

base dimension =???

OpenStudy (anonymous):

the base dimentions are just (5,5)

OpenStudy (anonymous):

not 5,5,5 (this includes the third dimension of height)

OpenStudy (anonymous):

its not 5,5

OpenStudy (anonymous):

i just tried that also...

OpenStudy (anonymous):

this is frustrating!

OpenStudy (anonymous):

check the units.. "in" for inches.. sometimes they ask for the units

OpenStudy (anonymous):

i guess i will just skip that one bc it is not working

OpenStudy (anonymous):

On a given day, the flow rate F (in cars per hour) on a congested roadway is given by the equation below, where v is the speed of traffic in miles per hour. F = v 22 + 0.07v2 What speed will maximize the flow rate on the road? (Round your answer to one decimal place.) mi/h

OpenStudy (anonymous):

do you know how to figure this out?

OpenStudy (anonymous):

Refer to a Mathematica solution attached.

OpenStudy (anonymous):

@robtobey thats cheating, lol

OpenStudy (anonymous):

and you did ignore the top surface